# Spacetime Quasicrystals

### v1

@@ IV.1.2 Symmetries of Lorentzian Lattices
  • For ${\rm I}_{s,1}$, the first $s$ fundamental roots form a sub-diagram of (sphereical/finite) type $B_{s}$, which corresponds to the fact that the odd Lorentzian lattices $\mathrm{I}_{s,1}$ are the Lorentzian analogues of the $B_{s}$ Euclidean root lattice ($\mathbb{Z}^{s}$).
@@ IV.1.3 Hyperbolic Coxeter Eigensystems, Salem Numbers, and Newman-Penrose Frames
If the group’s Coxeter diagram is bicolorable (i.e. if the vertices in the diagram can be colored red and blue, such that no two vertices of the same color are joined by an edge), then we can define $C_{{\rm red}}=\prod_{i\in{\rm red}}R_{i}$ as the product of all red fundamental reflections, $C_{{\rm blue}}=\prod_{i\in{\rm blue}}R_{i}$ as the product of all blue fundamental reflections, and the Coxeter element as $C=C_{{\rm red}}C_{{\rm blue}}$.
@@
    $\begin{array}[]{ccccccccc}V&=&V_{0}&\oplus&V_{1}&\oplus&\dots&\oplus&V_{k-1}\\ d&=&\mathbf{2}&\oplus&\mathbf{2}&\oplus&\dots&\oplus&\mathbf{2}\\ s,1&=&(1,1)&\oplus&(2,0)&\oplus&\dots&\oplus&(2,0).\end{array}$ (49)
@@ IV.2 Spacetime C&P Scheme
In the previous section, we first introduced the ordinary cut-and-project (“C&P”) scheme, followed by the more refined Symmetric C&P scheme, in which quasiperiodicity, orientational symmetry and discrete scale symmetry (inflation/deflation symmetry) are tightly interconnected. The most interesting quasicrystals are produced by the sC&P scheme. As discussed in Sec. III.2, an sC&P quasicrystal is determined by three choices, so in this section we review these three choices (in blue), and discuss any novelties associated with generalizing each one to the Lorentzian spacetime context.
@@ IV.3.1 Point Set, But No Tiling
So far have described the sC&P algorithm with symmetric window $\cal W$} as a way to generate a $G$-symmetric quasicrystalline point set (for some non-crystallographic group $G$). But in the Eucliean case (where $G$ is a finite group), one can often (as in the Penrose tiling, or Ammann-Beenker tiling) regard these points as the endpoints of edges; these edges as the boundaries of faces, etc, so that the $G$-symmetric point-set is promoted to a $G$-symmetric tiling or tessellation.
By contrast, in the Lorentzian case, one can’t promote a $G$-symmetric Lorentzian sC&P point set to a $G$-symmetric tiling (with a finite number of polygonal tiles). To see this, note that if the $G$-symmetric tiling contains any tile $T$, it must also contain the “rotated” tile gTgT ($\forall g\in G$); but if $G$ is an infinite group, this will either implies that the tile $T$ is accompanied by an infinite number of other, differently-oriented tiles; or else, $T$ must be invariant under a finite-index subgroup of $G$ (in which case the tile itself is “spherically symmetric”).
@@ IV.3.2 No Preferred Window Size
In the Euclidean case, if the window ${\cal W}\in V_{\rm in}$ is chosen to have a certain natural size, it uniquely lifts to a cell ${\cal C}\in V$ whose vertices have rational coordinates in the lattice basis, and whose $n$-faces project onto $V_{\rm ph}$ to give the $n$-tiles in the quasicrystalline tiling.
@@ IV.3.3 No Local Scale Invariance
In Section III.3.4, we explained two notions of scale invariance: local and global. And, in particular, in III.3.5 we explained that a C&P quasicrystal has local scale invariance – also known as “local inflation/deflation symmetry” (LIDS) – if and only if the inflated window $\widetilde{{\cal W}}$ can be obtained via a finite number of unions and intersections of translated copies of the original window $\cal W$}; and vice versa [7, 5]. For a spherically-symmetric window this is impossible (a rescaled ball cannot be expressed as a finite number of unions and intersections of translated copies of the original ball). So a $G$-symmetric spacetime quasicrystal (with infinite noncrystallographic symmetry group $G$) cannot have local inflation invariance.
@@ IV.4.1 Example 1: $(1+1)$-dimensional quasicrystals from $\mathrm{I}_{3,1}$
Now, in Fig. 11 we illustrate the resulting quasicrystal that results from several natural choices of the weighting function (or window): first, the self-dual quasicrystal obtained by choosing a Gaussian weighting function $W({\bf v}_{\rm in})$ with the self-dual width, i.e. $W({\bf v}_{\rm in})\propto{\rm exp}(-\pi v_{\rm in}^{2})$ with $v_{\rm in}^{2}\equiv\langle{\bf r}_{\rm in},{\bf r}_{\rm in}\rangle$ (left panel), the quasicrystal obtained by taking a circular-disk-shaped window $\cal W$} with radius $3\sigma$ (middle panel) where $\sigma=(2\pi)^{-1/2}$ is the standard deviation of the self-dual Gaussian, and the globally-scale-invariant quasicrystal of weight $w=1.01$ obtained by taking a weighting function $W({\bf v}_{\rm in})\propto v_{\rm in}^{-(2+1.01)}$.
@@ IV.4.2 Example 2: $(3+1)$-dimensional quasicrystals from $\mathrm{II}_{9,1}$
To construct a $(3+1)$ dimensional quasicrystal that is invariant under the group $G$ generated by $C$, we must choose the physical space to be spanned by the real pair of eigenvectors ${\bf v}_{0}^{\pm}$ and one of the four pairs of complex conjugate eigenvectors. For example, if we choose the first pair of complex conjugate eigenvectors ${\bf v}_{1}^{\pm}$ we obtain the projection operators
@@
Finally, for any of these four inequivalent decompositionsl, our final choice is the weighting function: once again, the two most natural choices are the Gaussian-self-dual weighting function, or the power-law weighting function which is globally scale invariant with weight $w$.
@@ V.1 The Universe in a Nutshell
To reconcile this with the $(3+1)$ dimensional universe we observe, the standard approach is to imagine (a modern variant of) the Kaluza-Klein picture in which four of the ten dimensions are large and non-compact, while the remaining six “extra” spatial dimensions are “compactified” (ı.e. curled up into an unobservably tiny 6D compact manifold). Alas, this picture so far does not have any observational evidence in its favour, and arguably creates more problems than its solves. In particular, stabilizing the six extra dimensions in a way that agrees with observations has been a notoriously difficult, and has driven theorists to consider awkwardly complex mechanisms that are often compared to Rube-Golberg machines, and lead to a vast “string landscape” of possible vacua for 4D spacetime.
@@ V.1.1 Classical Picture: Irrational Spacetime
Figure 12: Case 1: the line’s slope is a rational number $p/q$, with $p$ and $q$ relatively prime integers. Then, if one moves $q$ steps in the $x$-direction and $p$ steps in the $y$-direction, one retursn to the same point on the torus $\mathbb{R}^{2}/\mathbb{Z}^{2}$. Upon compactification (identification of the red edges on the square fundamental cell for the lattice), the origin becomes identified with any other lattice point. Thus, after compactification, the rational line will eventually come back to itself. However, if the line is irrationally sloped, then no point of the lattice will belong to the line. Such a line densely fills the torus and avoids crossing itself.
The most well studied compactifications of string theory are the toroidal compactifications: here one takes $m$ of the target-space dimensions to be curled up in an $m$-dimensional torus $T^{m}$, which is constructed by taking $m$-dimensional flat space $\mathbb{R}^{m}$ modulo an $m$-dimensional lattice $\Lambda_{m}$: $T^{m}=\mathbb{R}^{m}/\Lambda_{m}$. Due to stringy effects (in particular $T$-duality), such compactifications can have much higher symmetry than one would naively expect based on the isometry group of the compactified target spacetime. In Ref. [52], Moore studied such toroidal compactifications in generality, including the possibility of compatifying the time direction! He found that the most symmetric toroidal compactification of the superstring was obtained when all 10 dimensions of the target space were compactified on the 10D Lorentzian torus $T^{9,1}$ constructed by taking 10D Minkowski space $\mathbb{R}^{9,1}$ modulo the 10D even self-dual Lorentzian lattice $\Lambda_{9,1}$: $T^{9,1}=\mathbb{R}^{9,1}/\mathrm{II}_{9,1}$.
This most-symmetric superstring compactification was assumed to be unphysical, and of purely academic interest, for two reasons: (i) first, because if all ten dimensions are compactified at an ultra-microscopic scale, then one seems to be left without any remaining noncompact directions to serve as the $(3+1)$ macroscopic directions we observe; and (ii) second, because all directions (including time) are compactified, then the target spacetime has closed time-like curves which are usually assumed to be pathological, since they seem to be paradoxical and at odds with the causality properties of our universe.
@@
To see why, let us first warm up with a simplified 2D example: Consider the 2D square lattice $\Lambda_{2}$ in the 2D Euclidean plane $\mathbb{R}^{2}$; and a 1D line of slope $\sigma$ cutting through the lattice. Now if we mod out $\mathbb{R}^{2}$ by the 2D lattice $\Lambda_{2}$, we obtain the torus $T^{2}=\mathbb{R}^{2}/\Lambda_{2}$; and the 1D line in $\mathbb{R}^{2}$ is mapped to a 1D curve on the torus $T^{2}$ (see the blue line in Fig. 12). If the original 1D line in $\mathbb{R}^{2}$ had a rational slope $\sigma=p/q$ (where $p$ and $q$ are relatively prime integers), it is mapped to a closed curve of finite length on $T^{2}$, since the curve comes back joins up with itself after wrapping $p$ times around the $T^{2}$’s hole, and $q$ times through the $T^{2}$’s hole (top panel in Fig. 12). By contrast, if the original line in $\mathbb{R}^{2}$ had an irrational slope $\sigma$, it is mapped to an open curve of infinite length on $T^{2}$, since the curve never joins up with itself, and instead wraps around and around the $T^{2}$, densely covering it (bottom panel in Fig. 12). In this case, an ant confined to the line would inhabit a 1D world of infinite extent, even thought it is embedded in a compact 2D torus of finite area!
@@
Figure 13:  Upon compactification, all points along the blue line can be identified with a point within the first square, which has been turned into a 2D torus. By extension, the full infinite blue line is mapped within the square. The game resembles “Pac-Man”: whenever the line hits the boundary of the square, it restarts in the opposite edge from its equivalent point (up to edge-identification). This is depicted at the background within the zoomed-in square. Thus, the whole infinite 1D blue line can fit inside a tiny 2D torus. The no self-crossing condition follows from irrationality. Extrapolating to the idea that the blue line represents our 4D spacetime living within 10 dimensions, the whole infinite universe can fit in a tiny nutshell! Figure adapted from Hawking’s book [64].
@@
Now return to the 10D example of interest: Consider the 10D Lorentzian lattice $\mathrm{II}_{9,1}$ in 10D Minkowski space $\mathbb{R}^{9,1}$. We have seen in Sec. IV.4.2 that the Coxeter element $C$ of $\mathrm{II}_{9,1}$ generates a non-crystallographic group $G$ that group induces a decomposition $V=V_{\rm ph}\oplus V_{\rm in}$, where $V_{\rm ph}=\mathbb{R}^{3,1}$ (a copy of 4D Minkowski space) and $V_{\rm in}=\mathbb{R}^{6,0}$ (a copy of 6D Euclidean space) transform under two different representations ($\rho_{3,1}$ and $\rho_{6,0}$) of $G$. Crucially, since $G$ is non-crystallographic, this 4D Minkowksi subspace $V_{\rm ph}=\mathbb{R}^{3,1}$ is completely irrational with respect to the lattice $\mathrm{II}_{9,1}$ (i.e. it does not intersect any points in the lattice $\mathrm{II}_{9,1}$, except for the origin). So now, if we mod out $\mathbb{R}^{9,1}$ by the lattice $\mathrm{II}_{9,1}$, we obtain the 10D Lorentzian torus $T^{9,1}=\mathbb{R}^{9,1}/\mathrm{II}_{9,1}$; and the irrational 4D Minkowski subspace $V_{\rm ph}=\mathbb{R}^{3,1}$ is mapped to a (3+1)(3+1)D surface on the $T^{9,1}$, which densely fills the torus, wrapping around and around it without ever intersecting itself (like the blue line in the lower panel of Fig. 12).
@@
Wigner realized that different types of elementary particles are characterized by the way in which they transform under the fundamental symmetries of the theory. In particular, elementary particles that transfom as the representation $\rho_{3,1}$ of $G$ would be confined to the irrational $(3+1)$-dimensional surface that transforms in the appropriate way. A being made of such particles would regard themselves as living in an infinite (non-compact) 4D Minkowski space, with no closed timelike curves, despite the fact that this 4D spacetime is embedded in a 10D torus $T^{9,1}$ of finite volume, with all 10 dimensions compactified (including the time direction)!

### v2

@@ IV.1.2 Symmetries of Lorentzian Lattices
  • For ${\rm I}_{s,1}$, the first $s$ fundamental roots form a sub-diagram of (spherical/finite) type $B_{s}$, which corresponds to the fact that the odd Lorentzian lattices $\mathrm{I}_{s,1}$ are the Lorentzian analogues of the $B_{s}$ Euclidean root lattice, which is isomorphic to the hypercubic lattice $\mathbb{Z}^{s}$.
@@ IV.1.3 Hyperbolic Coxeter Eigensystems, Salem Numbers, and Newman-Penrose Frames
If the group’s Coxeter diagram is bicolorable (i.e. if the vertices in the diagram can be colored red and blue, such that no two vertices of the same color are joined by an edge; see Fig. 4), then we can define $C_{{\rm red}}=\prod_{i\in{\rm red}}R_{i}$ as the product of all red fundamental reflections, $C_{{\rm blue}}=\prod_{i\in{\rm blue}}R_{i}$ as the product of all blue fundamental reflections, and the Coxeter element as $C=C_{{\rm red}}C_{{\rm blue}}$.
@@
    $\begin{array}[]{ccccccccc}V&=&V_{0}&\oplus&V_{1}&\oplus&\dots&\oplus&V_{k-1}\\ d&=&\mathbf{2}&\oplus&\mathbf{2}&\oplus&\dots&\oplus&\mathbf{2}\\ (s,1)&=&(1,1)&\oplus&(2,0)&\oplus&\dots&\oplus&(2,0).\end{array}$ (49)
@@ IV.2 Spacetime C&P Scheme
In the previous section, we first introduced the ordinary cut-and-project (“C&P”) scheme, followed by the more refined Symmetric C&P scheme, in which quasi-periodicity, orientational symmetry and discrete scale symmetry (inflation/deflation symmetry) are tightly interconnected. The most interesting quasicrystals are produced by the sC&P scheme. As discussed in Sec. III.2, an sC&P quasicrystal is determined by three choices, so in this section we review these three choices (in blue), and discuss any novelties associated with generalizing each one to the Lorentzian spacetime context.
@@ IV.3.1 Point Set, But No Tiling
So far have described the sC&P algorithm with symmetric window $\cal W$} as a way to generate a $G$-symmetric quasi-crystalline point set (for some non-crystallographic group $G$). But in the Euclidean case (where $G$ is a finite group), one can often (as in the Penrose tiling, or Ammann-Beenker tiling) regard these points as the endpoints of edges; these edges as the boundaries of faces, etc, so that the $G$-symmetric point-set is promoted to a $G$-symmetric tiling or tessellation.
By contrast, in the Lorentzian case, one can’t promote a $G$-symmetric Lorentzian sC&P point set to a $G$-symmetric tiling (with a finite number of polygonal tiles). To see this, note that if the $G$-symmetric tiling contains any tile $T$, it must also contain the “rotated” tile gTgT ($\forall g\in G$); but if $G$ is an infinite group, this will either imply that the tile $T$ is accompanied by an infinite number of other, differently-oriented tiles; or else, $T$ must be invariant under a finite-index subgroup of $G$ (in which case the tile itself is “spherically symmetric”).
@@ IV.3.2 No Preferred Window Size
In the Euclidean case, if the window ${\cal W}\in V_{\rm in}$ is chosen to have a certain natural size, it uniquely lifts to a cell ${\cal C}\in V$ whose vertices have rational coordinates in the lattice basis, and whose $n$-faces project onto $V_{\rm ph}$ to give the $n$-tiles in the quasi-crystalline tiling.
@@ IV.3.3 No Local Scale Invariance
In Section III.3.4, we explained two notions of scale invariance: local and global. And, in particular, in III.3.5 we explained that a C&P quasicrystal has local scale invariance – also known as “local inflation/deflation symmetry” (LIDS) – if and only if the inflated window $\widetilde{{\cal W}}$ can be obtained via a finite number of unions and intersections of translated copies of the original window $\cal W$}; and vice versa [7, 5]. For a spherically-symmetric window this is impossible (a rescaled ball cannot be expressed as a finite number of unions and intersections of translated copies of the original ball). So a $G$-symmetric spacetime quasicrystal (with infinite non-crystallographic symmetry group $G$) cannot have local inflation invariance.
@@ IV.4.1 Example 1: $(1+1)$-dimensional Quasicrystals from $\mathrm{I}_{3,1}$
Now, in Fig. 11 we illustrate the resulting quasicrystal that results from several natural choices of the weighting function (or window): first, the self-dual quasicrystal obtained by choosing a Gaussian weighting function $W({\bf v}_{\rm in})$ with the self-dual width, i.e. $W({\bf v}_{\rm in})\propto{\rm exp}(-\pi v_{\rm in}^{2})$ with $v_{\rm in}^{2}\equiv\langle{\bf v}_{\rm in},{\bf v}_{\rm in}\rangle$ (top left panel), the quasicrystal obtained by taking a circular-disk-shaped window $\cal W$} with radius $3\sigma$ (top right panel) where $\sigma=(2\pi)^{-1/2}$ is the standard deviation of the self-dual Gaussian, and two globally-scale-invariant quasicrystals (bottom panel) of weight $k=1.01$ (left) and $k=0.25$ (right), obtained by taking a weighting function $W({\bf v}_{\rm in})\propto v_{\rm in}^{-(2+k)}$ for these two values of $k$.
@@ IV.4.2 Example 2: $(3+1)$-dimensional Quasicrystals from $\mathrm{II}_{9,1}$
To construct a $(3+1)$-dimensional quasicrystal that is invariant under the group $G$ generated by $C$, we must choose the physical space to be spanned by the real pair of eigenvectors ${\bf v}_{0}^{\pm}$ and one of the four pairs of complex conjugate eigenvectors. For example, if we choose the first pair of complex conjugate eigenvectors ${\bf v}_{1}^{\pm}$ we obtain the projection operators
@@
Finally, for any of these four inequivalent decompositions, our final choice is the weighting function: once again, the two most natural choices are the Gaussian-self-dual weighting function, or the power-law weighting function which is globally scale invariant with weight $w$.
@@ V.1 The Universe in a Nutshell
To reconcile this with the $(3+1)$ dimensional universe we observe, the standard approach is to imagine (a modern variant of) the Kaluza-Klein picture in which four of the ten dimensions are large and non-compact, while the remaining six “extra” spatial dimensions are “compactified” (ı.e. curled up into an unobservably tiny 6D compact manifold). Alas, this picture so far does not have any observational evidence in its favour, and arguably creates more problems than its solves. In particular, stabilizing the six extra dimensions in a way that agrees with observations has been a notoriously difficult, and has driven theorists to consider awkwardly complex mechanisms that are often compared to Rube-Goldberg machines, and lead to a vast “string landscape” of possible vacua for 4D spacetime.
@@ V.1.1 Classical Picture: Irrational Spacetime
Figure 12: Case I: the line’s slope is a rational number $p/q$, with $p$ and $q$ relatively prime integers. Then, if one moves $q$ steps in the $x$-direction and $p$ steps in the $y$-direction, one returns to the same point on the torus $\mathbb{R}^{2}/\mathbb{Z}^{2}$. Upon compactification (identification of the red edges on the square fundamental cell for the lattice), the origin becomes identified with any other lattice point. Thus, after compactification, the rational line will eventually come back to itself. However, if the line is irrationally sloped (case II), then no point of the lattice will belong to the line. Such a line densely fills the torus, extends indefinitely on it and avoids ever crossing itself as it wraps around the torus.
The most well studied compactifications of string theory are the toroidal compactifications: here one takes $m$ of the target-space dimensions to be curled up in an $m$-dimensional torus $T^{m}$, which is constructed by taking $m$-dimensional flat space $\mathbb{R}^{m}$ modulo an $m$-dimensional lattice $\Lambda_{m}$: $T^{m}=\mathbb{R}^{m}/\Lambda_{m}$. Due to stringy effects (in particular $T$-duality), such compactifications can have much higher symmetry than one would naively expect based on the isometry group of the compactified target spacetime. In Ref. [53], Moore studied such toroidal compactifications in generality, including the possibility of compactifying the time direction! He found that the most symmetric toroidal compactification of the superstring was obtained when all 10 dimensions of the target space were compactified on the 10D Lorentzian torus $T^{9,1}$ constructed by taking 10D Minkowski space $\mathbb{R}^{9,1}$ modulo the 10D even self-dual Lorentzian lattice $\Lambda_{9,1}$: $T^{9,1}=\mathbb{R}^{9,1}/\mathrm{II}_{9,1}$.
This most-symmetric superstring compactification was assumed to be unphysical, and of purely academic interest, for two reasons: (i) first, because if all ten dimensions are compactified at an ultra-microscopic scale, then one seems to be left without any remaining non-compact directions to serve as the $(3+1)$ macroscopic directions we observe; and (ii) second, because all directions (including time) are compactified, then the target spacetime has closed time-like curves which are usually assumed to be pathological, since they seem to be paradoxical and at odds with the causality properties of our universe.
@@
To see why, let us first warm up with a simplified 2D example: Consider the 2D square lattice $\Lambda_{2}$ in the 2D Euclidean plane $\mathbb{R}^{2}$; and a 1D line of slope $\sigma$ cutting through the lattice. Now if we mod out $\mathbb{R}^{2}$ by the 2D lattice $\Lambda_{2}$, we obtain the torus $T^{2}=\mathbb{R}^{2}/\Lambda_{2}$; and the 1D line in $\mathbb{R}^{2}$ is mapped to a 1D curve on the torus $T^{2}$ (see the blue line in Fig. 12). If the original 1D line in $\mathbb{R}^{2}$ had a rational slope $\sigma=p/q$ (where $p$ and $q$ are relatively prime integers), it is mapped to a closed curve of finite length on $T^{2}$, since the curve comes back joins up with itself after wrapping $p$ times around the $T^{2}$’s hole, and $q$ times through the $T^{2}$’s hole (top panel in Fig. 12). By contrast, if the original line in $\mathbb{R}^{2}$ had an irrational slope $\sigma$, it is mapped to an open curve of infinite length on $T^{2}$, since the curve never joins up with itself, and instead wraps around and around the $T^{2}$, densely covering it (bottom panel in Fig. 12). In this case, an ant confined to the line would inhabit a 1D world of infinite extent, even though it is embedded in a compact 2D torus of finite area!
@@
Figure 13:  Upon compactification, all points along the blue line can be identified with a point within the first square, which has been turned into a 2D torus. By extension, the full infinite blue line is mapped within the square. The game resembles “Pac-Man”: whenever the line hits the boundary of the square, it restarts in the opposite edge from its equivalent point (up to edge-identification). This is depicted at the background within the zoomed-in square. Thus, the whole infinite 1D blue line can fit inside a tiny 2D torus. The no self-crossing condition follows from irrationality. Extrapolating to the idea that the blue line represents our 4D spacetime living within 10 dimensions, the whole infinite universe can fit in a tiny nutshell! Figure adapted from Hawking’s [66].
@@
Now return to the 10D example of interest: consider the 10D Lorentzian lattice $\mathrm{II}_{9,1}$ in 10D Minkowski space $\mathbb{R}^{9,1}$. We have seen in Sec. IV.4.2 that the Coxeter element $C$ of $\mathrm{II}_{9,1}$ generates a non-crystallographic group $G$ that group induces a decomposition $V=V_{\rm ph}\oplus V_{\rm in}$, where $V_{\rm ph}=\mathbb{R}^{3,1}$ (a copy of 4D Minkowski space) and $V_{\rm in}=\mathbb{R}^{6,0}$ (a copy of 6D Euclidean space) transform under two different representations ($\rho_{3,1}$ and $\rho_{6,0}$) of $G$. Crucially, since $G$ is non-crystallographic, this 4D Minkowski subspace $V_{\rm ph}=\mathbb{R}^{3,1}$ is completely irrational with respect to the lattice $\mathrm{II}_{9,1}$ (i.e. it does not intersect any points in the lattice $\mathrm{II}_{9,1}$, except for the origin). So now, if we mod out $\mathbb{R}^{9,1}$ by the lattice $\mathrm{II}_{9,1}$, we obtain the 10D Lorentzian torus $T^{9,1}=\mathbb{R}^{9,1}/\mathrm{II}_{9,1}$; and the irrational 4D Minkowski subspace $V_{\rm ph}=\mathbb{R}^{3,1}$ is mapped to a (3+1)(3+1)D surface on the $T^{9,1}$, which densely fills the torus, wrapping around and around it without ever intersecting itself (like the blue line in the lower panel of Fig. 12).
@@
Wigner realized that different types of elementary particles are characterized by the way in which they transform under the fundamental symmetries of the theory. In particular, elementary particles that transform as the representation $\rho_{3,1}$ of $G$ would be confined to the irrational $(3+1)$-dimensional surface that transforms in the appropriate way. A being made of such particles would regard themselves as living in an infinite (non-compact) 4D Minkowski space, with no closed timelike curves, despite the fact that this 4D spacetime is embedded in a 10D torus $T^{9,1}$ of finite volume, with all 10 dimensions compactified (including the time direction)!